OpenStudy (anonymous):

A coin is flipped 150 times. The results of the experiment are shown in the following table: Head 63 Tail 87 Which of the following statements best describes the experimental probability of getting heads?

OpenStudy (anonymous):

It is equal to the theoretical probability. It is 8% lower than the theoretical probability. It is 8% higher than the theoretical probability. The experimental probability cannot be concluded from the data in the table.

OpenStudy (anonymous):

1/2*63/150?

OpenStudy (anonymous):

help

OpenStudy (mathstudent55):

What is the theoretical probability of getting heads in this case of 150 tosses?

OpenStudy (anonymous):

63

OpenStudy (mathstudent55):

No that is what happened in the experiment. In theory, if you toss a coin 150 times, how many heads do you expect?

OpenStudy (anonymous):

2

OpenStudy (mathstudent55):

?

OpenStudy (anonymous):

two side to a coin

OpenStudy (mathstudent55):

Each time you toss a coin, there is an equal probability of getting heads or tails. If you toss a coin 150 times, there is a theoretical probability of getting heads 75 times and tails 75 times. Do you understand my question now?

OpenStudy (anonymous):

yes i understand

OpenStudy (mathstudent55):

A coin is flipped 150 times Heads Tails Experimental outcome: 63 87 Theoretical outcome 75 75 Ok?

OpenStudy (mathstudent55):

Now let's look at percentages. The theoretical 75 heads are: 75/150 * 100 = 50% The actual heads are: 63/150 * 100 = 42% The actual tails are: 87/150 * 100 = 58%

OpenStudy (mathstudent55):

A coin is flipped 150 times Heads Tails Experimental outcome: 63 (42%) 87 (58%) Theoretical outcome 75 (50%) 75 (50%) Now look above at the "Heads" column. The theoretical heads result is 50%. The experimental (actual) heads result was 42%. Now read your choices and choose the correct answer.

OpenStudy (anonymous):

b?

OpenStudy (mathstudent55):

Exactly, since 42% is 8% less than 50%. Good job!

OpenStudy (anonymous):

@mathstudent55 thanks

OpenStudy (mathstudent55):

You're welcome.